# Clustering-based Low-Rank Matrix Approximation: An Adaptive Theoretical Analysis with Application to Data Compression

### v1

@@ Abstract
Low-rank matrix approximation (LoRMA) is a fundamental tool for compressing high-resolution data matrices by extracting important features while suppressing redundancy. Low-rank methods, such as global singular value decomposition (SVD), apply uniform compression across the entire data matrix, often ignoring important local variations and leading to the loss of fine structural details. To address these limitations, we introduce an adaptive LoRMA, which partitions data matrix into overlapping patches, groups structurally similar patches into several clusters using k-means, and performs SVD within each cluster. We derive the overall compression factor accounting for patch overlap and analyze how patch size influences compression efficiency and computational cost. While the proposed adaptive LoRMA method is applicable to any data exhibiting high local variation, we focus on medical imaging due to its pronounced local variability. We evaluate and compare our adaptive LoRMA against global SVD across four imaging modalities: MRI, ultrasound, CT scan, and chest X-ray. Results demonstrate that adaptive LoRMA effectively preserves structural integrity, edge details, and diagnostic relevance, as measured by peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), mean squared error (MSE), intersection over union (IoU), and edge preservation index (EPI). Adaptive LoRMA significantly minimizes block artifacts and residual errors, particularly in pathological regions, consistently outperforming global SVD in terms of PSNR, SSIM, IoU, EPI, and achieving lower MSE. Adaptive LoRMA prioritizes clinically salient regions while allowing aggressive compression in non-critical regions, optimizing storage efficiency. Although adaptive LoRMA requires higher processing time, its diagnostic fidelity justifies the overhead for high-compression applications.
Index Terms:
Low-rank matrix approximation, adaptive compression, cluster-based SVD, k-means clustering, medical imaging
I Introduction
@@
In many modern data-processing applications, one of the primary challenges is dealing with high-resolution and large-scale data that contains significant redundancy [1, 2, 3]. Low-rank matrix approximation (LoRMA) has become a fundamental tool in machine learning and other domains due to its ability to capture essential structures in high-resolution and large-scale data while suppressing noise and redundancy. By representing a high-dimensional data matrix with a lower-rank approximation, LoRMA reduces storage requirements and computational costs while preserving essential structural information. A widely used approach to LoRMA is the singular value decomposition (SVD) [4], which provides an optimal low-rank approximation under the Frobenius norm, as demonstrated by the Eckart-Young theorem [5]. However, despite its effectiveness in many domains, the global nature of truncated SVD imposes limitations when applied to real-world data, particularly when local variations are critical. In practice, many data matrices exhibit gradual decay in their singular values, making it challenging to choose a hard threshold for truncation without risking the loss of important features [6, 7, 8]. Truncated SVD (what we refer in this paper as global SVD) applies uniform compression across the entire data matrix, disregarding local variations in structural complexity. This uniformity assumes homogeneity in the data’s structure, an assumption that rarely holds true for many data matrices [9, 10, 11]. For example, while some regions of a data matrix may be relatively smooth, others can contain important features. The global SVD technique not only fails to adapt to these local variations but can also introduce over-smoothing and block artifacts, leading to the loss of important features [12, 7, 13, 14, 15]. Furthermore, the combinatorial nature of rank minimization makes strict rank constraints computationally intractable for large-scale data matrices[16, 17].
@@
Local low-rank matrix approximation (LLoRMA) was proposed to address these limitations. The LLoRMA technique, initially introduced by Lee [18] in 2013, was applied to recommendation systems to improve prediction accuracy by capturing local structures within data matrices. Their experiments utilized datasets such as MovieLens, EachMovie, and Netflix, focusing on improving user preference predictions. While their primary focus was on recommendation systems, subsequent research has extended the application of LLoRMA to other domains, such as hyperspectral imaging [19, 20], dynamic MRI reconstruction [21, 22], and so on. Other studies, such as non-local means denoising [23, 24, 25, 26, 27] and block-matching 3D filtering [28, 29, 30, 31], demonstrated the benefits of patch-based processing for preserving and extracting important image features. Building on these ideas, this paper proposes a cluster-based SVD mathematical framework. The cluster-based SVD technique partitions a data matrix into overlapping patches, groups similar patches using k-means, and applies SVD within each cluster. By adapting the compression factor based on the local data matrix features, the method aims to preserve and extract important features while reducing storage burdens.
@@
The rest of this paper is organized as follows: Section II introduces low-rank matrix approximation and its challenges; Section III describes our proposed cluster-based SVD method in detail. Section IV introduces the evaluation metrics used to assess the performance of our proposed cluster-based SVD method. Section V discusses the experimental setup and the results. Section VI discusses the findings and limitations, and Section VII concludes with key insights and future directions.
II Technical background
II-A Low-rank matrix approximation
Many real-world data matrices exhibit redundancy [3], meaning their information can be represented using a lower-dimensional subspace without significant loss. This property motivates the use of low-rank approximation, where a data matrix is approximated by another data matrix of reduced rank, preserving essential structures and extracting important features while discarding noise and redundant information. A fundamental problem in low-rank approximation is to find a rank-$r_{g}$ matrix $\mathbf{X}$ that best approximates a given data matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ under the Frobenius norm
@@ II-A Low-rank matrix approximation
II-B Global SVD
II-B1 Theoretical overview
SVD decomposes a data matrix, $\mathbf{A}\in\mathbb{R}^{m\times n}$ ($m>n$), into three independent matrices:
@@ II-B1 Theoretical overview
where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are unitary matrices, $\mathbf{\Sigma}$ is a diagonal matrix and T denotes the transpose. The diagonal entries $\sigma_{i}$ of $\mathbf{\Sigma}$ are uniquely determined by $\mathbf{A}$ and are known as the singular values of $\mathbf{A}$. The singular values are arranged in descending order, that is, $\sigma_{1}\geq\sigma_{2}\geq\sigma_{3}\cdots\geq\sigma_{n}$. The columns of $\mathbf{U}$ and the columns of $\mathbf{V}$ are the left and right singular vectors of $\mathbf{A}$ respectively. The vectors $\mathbf{u}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{n}$ are the orthonormal eigenvectors of $\mathbf{AA}^{T}$ and $\mathbf{A}^{T}\mathbf{A}$ respectively. By retaining only $r_{g}$ (where $1<r_{g}<n$) singular values, we obtain a compressed data matrix, $\mathbf{A}_{r_{g}}$, which is an approximation of $\mathbf{A}$. That is,
@@
where $\mathbf{U}_{r_{g}}$ denotes a matrix of size $m\times r_{g}$ containing the first $r_{g}$ columns of $\mathbf{U}$, $\mathbf{\Sigma}_{r_{g}}$ denotes a diagonal matrix of size $r_{g}\times r_{g}$ containing $r_{g}$ singular values, and $\mathbf{V}_{r_{g}}^{T}$ denotes a matrix of size $n\times r_{g}$ containing the first $r_{g}$ columns of $\mathbf{V}$. Each value of $\sigma_{i}$ quantifies how much the corresponding component $\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$ contributes in the reconstruction of $\mathbf{A}$, that is, the larger the value of $\sigma_{i}$ the more $\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$ contributes to the reconstruction of $\mathbf{A}$. The computational cost for reconstructing $\mathbf{A}$ is approximately given by
@@
II-B2 Data matrix compression
@@
[…]
@@
III Method
III-A Overall workflow
The proposed method exploits the local and non-local self-similarity of data matrices by first grouping similar patches using kmeans and then applying SVD within each cluster. This enables adaptive compression while preserving and extracting important features. In contrast to global SVD (shown in Fig. 1(a)), which compresses the entire matrix uniformly, our approach enables localized compression based on structural complexity. Fig. 1(b) illustrates the main steps of our proposed cluster-based SVD method. The process starts with patch extraction, where the input data matrix is divided into overlapping patches. Next, the patches are grouped using k-means clustering to organize patches with similar structural patterns into different clusters. Each cluster is then processed independently with truncated SVD, where the number of retained singular values retained is adapted to the local complexity of the cluster. The compressed cluster matrices are then aggregated and reconstructed to obtain the final compressed matrix. This ensures that only the most informative components are retained, while redundant and less important pixels are removed.
Figure 1: Overview of the cluster-based and global SVD for data compression. In cluster-based SVD, the original data matrix $\mathbf{A}$ is first split into overlapping $p\times p$ patches $\{\mathbf{P}_{i}\}_{i=1}^{N}$. Each patch is vectorized into $\mathbb{R}^{p^{2}}$ and grouped into $K$ clusters $\{\mathcal{C}_{k}\}_{k=1}^{K}$ using k‑means, which are used to form cluster matrices $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$. SVD is then applied independently to each $\mathbf{C}_{k}$ to produce compressed cluster matrices $\widehat{\mathbf{C}}_{k}$. Finally, the compressed patches are aggregated to reconstruct the compressed data matrix $\widehat{\mathbf{A}}$. While global SVD applies truncated SVD directly to the full matrix $\mathbf{A}$, yielding low-rank components $\mathbf{U}_{r_{g}}$, $\boldsymbol{\Sigma}_{r_{g}}$, and $\mathbf{V}_{r_{g}}^{T}$, from which the compressed matrix $\mathbf{A}_{r_{g}}$ is reconstructed.
III-B Patch grouping
In this paper, we use the k-means clustering algorithm to group similar patches, thereby exploiting the local and non-local self-similarity of the data matrices [45, 46]. The idea is to identify and group patches that have similar features, even if they are spatially distant from each other. The k-means algorithm is used for proof of concept in this paper and other similar measure methods could also be considered; however this is not the aim for this paper, and we leave it for future work. To group similar patches using k-means clustering, we divide the data matrix into overlapping patches (this helps to capture more contextual information and reduce the boundary artifacts) each of size $p\times p$ (as shown in Algorithm 1, line 5). Each patch $\boldsymbol{P}_{k}$ is vectorized, that is,
@@ III-B Patch grouping
[…]
@@
Algorithm 1 describes how to extract overlapping patches from a data matrix and group similar patches using k-means clustering. It iteratively assigns patches to the nearest centroid and updates centroids until convergence.
@@
1:Input: Data matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, patch size $p$, number of clusters $K$
@@
12:         Compute distances: $d(j)=\|\mathbf{p}_{k}-\boldsymbol{\mu}_{j}\|_{2}$ for $j=1,\dots,K$triangleright $d(j)$ is the Euclidean distance between patch $\mathbf{p}_{k}$ and centroid $\boldsymbol{\mu}_{j}$
@@
III-C Cluster-dependent SVD
Once the patches have been grouped into clusters, each cluster is represented as a matrix. For a cluster $\mathcal{C}_{k}$ containing $N_{\mathbf{C}_{k}}$ patches, each represented by the vector $\mathbf{p}_{i}\in\mathbb{R}^{p^{2}}$, the corresponding matrix representation is given by
$\mathbf{C}_{k}=\left(\begin{array}[]{cccc}\mathbf{p}_{1}&\mathbf{p}_{2}&\cdots&\mathbf{p}_{N_{\mathbf{C}_{k}}}\end{array}\right)^{T}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}},$ (29)
@@ III-C Cluster-dependent SVD
$\Gamma(\mathbf{X})=\{\mathbf{U},\mathbf{\Sigma},\mathbf{V}\}\quad\text{such that}\quad\mathbf{X}=\mathbf{U}\,\mathbf{\Sigma}\,\mathbf{V}^{T}.$ (32)
@@
$CF_{\mathbf{C}_{i}}(r_{\mathbf{C}_{i}},N_{\mathbf{C}_{i}},p)\gg CF_{\mathbf{C}_{j}}(r_{\mathbf{C}_{j}},N_{\mathbf{C}_{j}},p).$ (35)
@@
Fig. 3 shows the intra-cluster variability, measured using the Frobenius norm, across different number of clusters $K$. The x-axis represents the number of clusters, while the y-axis quantifies the intra-cluster variability. The results demonstrates that when $K$ is small, patches with different features are grouped in the same cluster $\mathcal{C}_{k}$ resulting in cluster matrices $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$ that exhibit high variability. The variability within $\mathbf{C}_{k}$ limits the effectiveness of the low-rank approximation as the SVD of $\mathbf{C}_{k}$ cannot efficiently isolate meaningful low-rank structures. In particular, $\|\mathbf{C}_{k}-\widehat{\mathbf{C}}_{k}\|_{F}^{2}$ remains large for small $r_{\mathbf{C}_{k}}$, which leads to poor reconstruction fidelity. As a result, local matrix details and structural information are insufficiently captured, leading to artifacts or blurring in the reconstructed matrix. Conversely, as $K$ increases the patches within each cluster become more homogeneous reducing the variability in $\mathbf{C}_{k}$. This makes it easier to maintain fewer singular values $r_{\mathbf{C}_{k}}$ while still achieving accurate approximations. This is reflected in the decay of the singular values, where for more homogeneous matrices the tail singular values $\{\sigma_{\mathbf{C}_{k},i}\}_{i=r_{\mathbf{C}_{i}}+1}^{\min\{N_{\mathbf{C}_{k}},p^{2}\}}$ are smaller, reducing the reconstruction error. A larger $K$ leads to better data matrix quality, as the local and non-local self-similarity of the patches is effectively utilized.
@@
III-D The overall compression factor
The compression factor of the data matrix $\mathbf{C}_{k}$ for cluster $\mathcal{C}_{k}$ is determined by the ratio between the number of elements in the original matrix and the number of elements retained after applying truncated SVD. The data matrix $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$ contains $N_{\mathbf{C}_{k}}\times p^{2}$ elements before compression. After applying SVD and retaining only the top $r_{\mathbf{C}_{k}}$ singular values, the number of elements in the compressed representation is reduced to $r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1)$, which includes $r_{\mathbf{C}_{k}}$ singular values, $r_{\mathbf{C}_{k}}\times N_{\mathbf{C}_{k}}$ elements from the left singular matrix $\mathbf{U}_{r_{\mathbf{C}_{k}}}$, and $r_{\mathbf{C}_{k}}\times p^{2}$ elements from the right singular matrix $\mathbf{V}_{r_{\mathbf{C}_{k}}}$. Therefore if the original matrix is split into non-overlapping patches then the compression factor for cluster $\mathcal{C}_{k}$ is given by
$CF_{\mathbf{C}_{k}}(r_{\mathbf{C}_{k}},N_{\mathbf{C}_{k}},p)=\frac{N_{\mathbf{C}_{k}}p^{2}}{r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1)},$ (38)

### v2

@@ Abstract
Medical images are inherently high-resolution and contain locally varying structures that are crucial for diagnosis. Efficient compression of such data must therefore preserve diagnostic fidelity while minimizing redundancy. Low-rank matrix approximation (LoRMA) techniques have shown strong potential for image compression by capturing global correlations; however, they often fail to adapt to local structural variations across regions of interest, i.e., tumor areas or other diagnostically significant regions. To address these limitations, we introduce an adaptive LoRMA, which partitions a medical image into overlapping patches, groups structurally similar patches into several clusters using k-means, and performs SVD within each cluster. We derive the overall compression factor accounting for patch overlap and analyze how patch size influences compression efficiency and computational cost. While the proposed adaptive LoRMA method is applicable to any data exhibiting high local variation, we focus on medical imaging due to its pronounced local variability. We evaluate and compare our adaptive LoRMA against global SVD across four imaging modalities: MRI, ultrasound, CT scan, and chest X-ray. Results demonstrate that adaptive LoRMA effectively preserves structural integrity, edge details, and diagnostic relevance, as measured by peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), mean squared error (MSE), intersection over union (IoU), and edge preservation index (EPI). Adaptive LoRMA significantly minimizes block artifacts and residual errors, particularly in pathological regions, consistently outperforming global SVD in terms of PSNR, SSIM, IoU, EPI, and achieving lower MSE. Adaptive LoRMA prioritizes clinically salient regions while allowing aggressive compression in non-critical regions, optimizing storage efficiency. Although adaptive LoRMA requires higher processing time, its diagnostic fidelity justifies the overhead for high-compression applications.
keywords:
[…]
1 Introduction
@@
Modern medical imaging modalities such as MRI, CT, ultrasound, and X-ray produce high-resolution images that enable precise clinical diagnosis but pose significant challenges in terms of storage, transmission, and computational efficiency [40, 60, 50]. Low-rank matrix approximation (LoRMA) has become a fundamental tool in machine learning and other domains due to its ability to capture essential structures in high-resolution and large-scale data while suppressing noise and redundancy. By representing a high-resolution medical image with a lower-rank approximation, LoRMA reduces storage requirements and computational costs while preserving essential structural information. A widely used approach to LoRMA is the singular value decomposition (SVD) [35], which provides an optimal low-rank approximation under the Frobenius norm, as demonstrated by the Eckart-Young theorem [22]. However, despite its effectiveness in many domains, the global nature of truncated SVD imposes limitations when applied to real-world data, particularly when local variations are critical. In medical imaging, for example, many medical images exhibit gradual decay in their singular values, making it challenging to choose a hard threshold for truncation without risking the loss of important features [30, 36, 14]. Truncated SVD (what we refer in this paper as global SVD) applies uniform compression across the entire medical image, disregarding local variations in structural complexity. This uniformity assumes homogeneity in the data’s structure, an assumption that rarely holds true for many medical images [51, 9, 29]. For example, while some regions of a medical image may be relatively smooth, others can contain important features. The global SVD technique not only fails to adapt to these local variations but can also introduce over-smoothing and block artifacts, leading to the loss of important features [27, 36, 26, 28, 1]. Furthermore, the combinatorial nature of rank minimization makes strict rank constraints computationally intractable for large-scale data matrices [49, 13].
@@
Local low-rank matrix approximation (LLoRMA) was proposed to address these limitations. The LLoRMA technique, initially introduced by Lee [37] in 2013, was applied to recommendation systems to improve prediction accuracy by capturing local structures within data matrices. Their experiments utilized datasets such as MovieLens, EachMovie, and Netflix, focusing on improving user preference predictions. While their primary focus was on recommendation systems, subsequent research has extended the application of LLoRMA to other domains, such as hyperspectral imaging [54, 31], dynamic MRI reconstruction [48, 58], and so on. Other studies, such as non-local means denoising [16, 44, 45, 11, 21] and block-matching 3D filtering [17, 15, 25, 24], demonstrated the benefits of patch-based processing for preserving and extracting important image features. Building on these ideas, this paper proposes a cluster-based SVD mathematical framework. The cluster-based SVD technique partitions a medical image into overlapping patches, groups similar patches using k-means, and applies SVD within each cluster. By adapting the compression factor based on the local medical image features, the method aims to preserve and extract important features while reducing storage burdens.
@@
The rest of this paper is organized as follows: Section 2 introduces low-rank matrix approximation and its challenges; Section 3 describes our proposed cluster-based SVD method in detail. Section 4 introduces the evaluation metrics used to assess the performance of our proposed cluster-based SVD method. Section 5 discusses the experimental setup and the results. Section 6 discusses the findings and limitations, and Section 7 concludes with key insights and future directions.
2 Technical background
2.1 Low-rank matrix approximation
Many real-world data matrices exhibit redundancy [50], meaning their information can be represented using a lower-dimensional subspace without significant loss. This property motivates the use of low-rank approximation, where a medical image is approximated by another medical image of reduced rank, preserving essential structures and extracting important features while discarding noise and redundant information. A fundamental problem in low-rank approximation is to find a rank-$r_{g}$ matrix $\mathbf{X}$ that best approximates a given medical image $\mathbf{A}\in\mathbb{R}^{m\times n}$ under the Frobenius norm
@@ 2.1 Low-rank matrix approximation
2.2 Global SVD
2.2.1 Theoretical overview
SVD decomposes a medical image, $\mathbf{A}\in\mathbb{R}^{m\times n}$ ($m>n$), into three independent matrices:
@@ 2.2.1 Theoretical overview
where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are unitary matrices, $\mathbf{\Sigma}$ is a diagonal matrix and T denotes the transpose. The diagonal entries $\sigma_{i}$ of $\mathbf{\Sigma}$ are uniquely determined by $\mathbf{A}$ and are known as the singular values of $\mathbf{A}$. The singular values are arranged in descending order, that is, $\sigma_{1}\geq\sigma_{2}\geq\sigma_{3}\cdots\geq\sigma_{n}$. The columns of $\mathbf{U}$ and the columns of $\mathbf{V}$ are the left and right singular vectors of $\mathbf{A}$ respectively. The vectors $\mathbf{u}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{n}$ are the orthonormal eigenvectors of $\mathbf{AA}^{T}$ and $\mathbf{A}^{T}\mathbf{A}$ respectively. By retaining only $r_{g}$ (where $1<r_{g}<n$) singular values, we obtain a compressed medical image, $\mathbf{A}_{r_{g}}$, which is an approximation of $\mathbf{A}$. That is,
@@
where $\mathbf{U}_{r_{g}}$ denotes a matrix of size $m\times r_{g}$ containing the first $r_{g}$ columns of $\mathbf{U}$, $\mathbf{\Sigma}_{r_{g}}$ denotes a diagonal matrix of size $r_{g}\times r_{g}$ containing $r_{g}$ singular values, and $\mathbf{V}_{r_{g}}^{T}$ denotes a matrix of size $r_{g}\times n$ containing the first $r_{g}$ columns of $\mathbf{V}$. Each value of $\sigma_{i}$ quantifies how much the corresponding component $\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$ contributes in the reconstruction of $\mathbf{A}$, that is, the larger the value of $\sigma_{i}$ the more $\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}$ contributes to the reconstruction of $\mathbf{A}$. The computational cost for reconstructing $\mathbf{A}$ is approximately given by
@@
2.2.2 Data matrix compression
@@
where $\lceil\cdot\rceil$ gives the smallest integer. The value of $r_{g}$ directly influences the degree of compression achieved for a medical image. If fewer singular values are retained, more signal is discarded, resulting in higher compression but possibly lower matrix quality. Conversely, retaining more singular values preserves more signal, resulting in higher approximated matrix quality but less compression. The compression loss can be computed as
@@
3 Method
3.1 Overall workflow
The proposed method exploits the local and non-local self-similarity of medical images by first grouping similar patches using k-means and then applying SVD within each cluster. This enables adaptive compression while preserving and extracting important features. In contrast to global SVD (shown in Fig. 1(a)), which compresses the entire matrix uniformly, our approach enables localized compression based on structural complexity. Fig. 1(b) illustrates the main steps of our proposed cluster-based SVD method. The process starts with patch extraction, where the input medical image is divided into overlapping patches. Next, the patches are grouped using k-means clustering to organize patches with similar structural patterns into different clusters. Each cluster is then processed independently with truncated SVD, where the number of singular values retained is adapted to the local complexity of the cluster. The compressed cluster matrices are then aggregated and reconstructed to obtain the final compressed matrix. This ensures that only the most informative components are retained, while redundant and less important pixels are removed.
Figure 1: Overview of the cluster-based and global SVD for data compression. (a) Global SVD applies truncated SVD directly to the original medical image $\mathbf{A}$, yielding low-rank components $\mathbf{U}_{r_{g}}$, $\boldsymbol{\Sigma}_{r_{g}}$, and $\mathbf{V}_{r_{g}}^{T}$, from which the compressed medical image $\mathbf{A}_{r_{g}}$ is reconstructed. (b) In cluster-based SVD, the original medical image $\mathbf{A}$ is first split into overlapping $p\times p$ patches $\{\mathbf{P}_{i}\}_{i=1}^{N}$. Each patch is vectorized into $\mathbb{R}^{p^{2}}$ and grouped into $K$ clusters $\{\mathcal{C}_{k}\}_{k=1}^{K}$ using k‑means, which are used to form cluster matrices $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$. SVD is then applied independently to each $\mathbf{C}_{k}$ to produce compressed cluster matrices $\widehat{\mathbf{C}}_{k}$. Finally, the compressed patches are aggregated to reconstruct the compressed medical image $\widehat{\mathbf{A}}$.
3.2 Patch grouping
In this paper, we use the k-means clustering algorithm to group similar patches, thereby exploiting the local and non-local self-similarity of the medical images [10, 18]. The idea is to identify and group patches that have similar features, even if they are spatially distant from each other. The k-means algorithm is used for proof of concept in this paper and other similar measure methods could also be considered; however this is not the aim for this paper, and we leave it for future work. To group similar patches using k-means clustering, we divide the medical image into overlapping patches (this helps to capture more contextual information and reduce the boundary artifacts) each of size $p\times p$ (as shown in Algorithm 1, line 5). Each patch $\boldsymbol{P}_{k}$ is vectorized, that is,
@@ 3.2 Patch grouping
where $\langle\cdot\rangle$ is the inner product. It then follows that
@@
Algorithm 1 describes how to extract overlapping patches from a medical image and group similar patches using k-means clustering. It iteratively assigns patches to the nearest centroid and updates centroids until convergence.
@@
1:Input: medical image $\mathbf{A}\in\mathbb{R}^{m\times n}$, patch size $p$, number of clusters $K$
@@
12:   Compute distances: $d(j)=\|\mathbf{p}_{k}-\boldsymbol{\mu}_{j}\|_{2}$ for $j=1,\dots,K$ $\triangleright$ $d(j)$ is the Euclidean distance between patch $\mathbf{p}_{k}$ and centroid $\boldsymbol{\mu}_{j}$
@@
3.3 Cluster-dependent SVD
Once the patches have been grouped into clusters, each cluster is represented as a matrix. For a cluster $\mathcal{C}_{k}$ containing $N_{\mathbf{C}_{k}}$ patches, each represented by the vector $\mathbf{p}_{i}\in\mathbb{R}^{p^{2}}$, the corresponding medical image representation is given by
$\mathbf{C}_{k}=\left(\begin{array}[]{cccc}\mathbf{p}_{1}^{T}&\mathbf{p}_{2}^{T}&\cdots&\mathbf{p}_{N_{\mathbf{C}_{k}}}^{T}\end{array}\right)^{T}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}},$ (29)
@@ 3.3 Cluster-dependent SVD
$\Gamma(\mathbf{X})=\{\mathbf{U}_{r_{g}},\mathbf{\Sigma}_{r_{g}},\mathbf{V}_{r_{g}}\}\quad\text{such that}\quad\mathbf{X}=\mathbf{U}_{r_{g}}\,\mathbf{\Sigma}_{r_{g}}\,\mathbf{V}_{r_{g}}^{T}.$ (32)
@@
$CF_{\mathbf{C}_{i}}(p,N_{\mathbf{C}_{j}},r_{\mathbf{C}_{j}})\gg CF_{\mathbf{C}_{j}}(p,N_{\mathbf{C}_{j}},r_{\mathbf{C}_{j}}).$ (35)
@@
Fig. 3 shows the intra-cluster variability, measured using the Frobenius norm, across different number of clusters $K$. The x-axis represents the number of clusters, while the y-axis quantifies the intra-cluster variability. The results demonstrates that when $K$ is small, patches with different features are grouped in the same cluster $\mathcal{C}_{k}$ resulting in cluster matrices $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$ that exhibit high variability. The variability within $\mathbf{C}_{k}$ limits the effectiveness of the low-rank approximation as the SVD of $\mathbf{C}_{k}$ cannot efficiently isolate meaningful low-rank structures. In particular, $\|\mathbf{C}_{k}-\widehat{\mathbf{C}}_{k}\|_{F}^{2}$ remains large for small $r_{\mathbf{C}_{k}}$, which leads to poor reconstruction fidelity. As a result, local matrix details and structural information are insufficiently captured, leading to artifacts or blurring in the reconstructed matrix. Conversely, as $K$ increases the patches within each cluster become more homogeneous reducing the variability in $\mathbf{C}_{k}$. This makes it easier to maintain fewer singular values $r_{\mathbf{C}_{k}}$ while still achieving accurate approximations. This is reflected in the decay of the singular values, where for more homogeneous matrices the tail singular values $\{\sigma_{\mathbf{C}_{k},i}\}_{i=r_{\mathbf{C}_{i}}+1}^{\min\{N_{\mathbf{C}_{k}},p^{2}\}}$ are smaller, reducing the reconstruction error. A larger $K$ leads to better medical image quality, as the local and non-local self-similarity of the patches is effectively utilized.
@@
3.4 The overall compression factor
The compression factor of the medical image $\mathbf{C}_{k}$ for cluster $\mathcal{C}_{k}$ is determined by the ratio between the number of elements in the original matrix and the number of elements retained after applying truncated SVD. The medical image $\mathbf{C}_{k}\in\mathbb{R}^{N_{\mathbf{C}_{k}}\times p^{2}}$ contains $N_{\mathbf{C}_{k}}\times p^{2}$ elements before compression. After applying SVD and retaining only the top $r_{\mathbf{C}_{k}}$ singular values, the number of elements in the compressed representation is reduced to $r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1)$, which includes $r_{\mathbf{C}_{k}}$ singular values, $r_{\mathbf{C}_{k}}\times N_{\mathbf{C}_{k}}$ elements from the left singular matrix $\mathbf{U}_{r_{\mathbf{C}_{k}}}$, and $r_{\mathbf{C}_{k}}\times p^{2}$ elements from the right singular matrix $\mathbf{V}_{r_{\mathbf{C}_{k}}}$. Therefore if the original matrix is split into non-overlapping patches then the compression factor for cluster $\mathcal{C}_{k}$ is given by
$CF_{\mathbf{C}_{k}}(p,N_{\mathbf{C}_{k}},r_{\mathbf{C}_{k}})=\frac{N_{\mathbf{C}_{k}}p^{2}}{r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1)},$ (38)